Hello. I work for a veterinarian that likes to mix solutions using injectable medications. He likes to make diluted solutions as he uses very small doses. I will give an example as I understand what he wants but cannot fully understand the mathematics behind making the solution.
I have a medication X that is concentrated at 25mg/ml.
I want to make up a solution to dilute this to 0.05mg/ml with a target volume of 4ml.
How much of medication X needs to be mixed with how much sterile saline (which will just dilute it, that's all) to make the total of 4ml.
It is the math to figure out how much of medication X that I need. Also, I cannot figure out if the target volume has any effect on the calculations.
Please help! I want to understand this!
Thank you.

Answer by Alan3354(69443)

I have a medication X that is concentrated at 25mg/ml.
I want to make up a solution to dilute this to 0.05mg/ml with a target volume of 4ml.
How much of medication X needs to be mixed with how much sterile saline (which will just dilute it, that's all) to make the total of 4ml.
It is the math to figure out how much of medication X that I need. Also, I cannot figure out if the target volume has any effect on the calculations.
=========================
I have a medication X that is concentrated at 25mg/ml.
I want to make up a solution to dilute this to 0.05mg/ml with a target volume of 4ml.
How much of medication X needs to be mixed with how much sterile saline (which will just dilute it, that's all) to make the total of 4ml.
t = amount of 25 mg/ml
s = amount of saline
---
t + s = 4 (total solution)
25t + 0*s = 0.05*(t+s) (total medication)
---
t + s = 4 --> s = 4-t
25t + 0*s = 0.05*(t+s)
Sub for s
25t = 0.05*(t + 4-t)
25t = 0.05*4 = 0.2
t = 0.2/25
t = 0.008 ml of the 25 mg/ml
----------------
That's how mixing 2 solutions is done.
Straight dilution is simpler. Diluting from 25 to 0.05 is 1/500.
For 4 ml of 0.05 you need 4/500 ml of the original = 0.008 ml
----------------------
That's not much, might be hard to measure.
It takes very little because it's being diluted to 1/500 of its concentration.